# Stopping Rules for Stochastic Gradient Descent via Anytime-Valid Confidence Sequences

### v3

@@ 3 Anytime-Valid Stopping Rules
Thus $\mathcal{E}_{\alpha}\subseteq\{f(\bar{x}_{\tau_{\varepsilon}})-f^{\star}\leq\varepsilon\}$, and taking probabilities yields
$\mathbb{P}\Big(f(\bar{x}_{\tau_{\varepsilon}})-f^{\star}\leq\varepsilon\Big)\;\geq\;\mathbb{P}(\mathcal{E}_{\alpha})\;\geq\;1-\alpha,$
which is exactly (33). This proves part (ii) and completes the proof. ∎
Theorem 3.1 shows that projected SGD equipped with the stopping rule $\tau_{\varepsilon}$ is a data-dependent $\varepsilon$-optimal procedure: one may monitor $U_{t}^{\mathrm{obs}}(\alpha)$ online and stop as soon as it falls below $\varepsilon$, without specifying a horizon in advance or invalidating the $1-\alpha$ guarantee. This yields fully implementable stopping rules based only on running sums of $\eta_{t}$, $\eta_{t}^{2}$, and the observed gradient norms, with the usual tradeoff between adaptivity and reliance on global problem parameters.
In the next subsection we study the stopping-time complexity associated with the rule $U_{t}^{\mathrm{obs}}(\alpha)\leq\varepsilon$. We present explicit bounds with deterministic constants via a strengthening of our assumptions on the stochastic gradients.
3.1 Stopping-time complexity
In this subsection we derive explicit bounds on the stopping time $\tau_{\varepsilon}$ under standard stepsize schedules for stochastic approximation. Our assumptions on the stochastic gradients are as follows.
Assumption 3.2 (Stochastic gradients (strengthened)).
For all $t\geq 1$, the stochastic gradient $g_{t}$ satisfies:
  • (i)
  • (ii)
    Boundedness: $\|g_{t}\|\leq G$ almost surely.
Note that Assumption 3.2(ii) implies Assumption 1.1(ii), and moreover implies the conditional sub-Gaussian Assumption 1.1(iii) with variance proxy $\sigma^{2}:=4G^{2}$. Indeed, for any $\lambda\in\mathbb{R}$ and $u\in\mathbb{R}^{d}$,
$|\big\langle g_{t}-\mathbb{E}[g_{t}\mid\mathcal{F}_{t-1}],u\big\rangle|\leq 2G\|u\|\qquad\text{almost surely},$
so by Hoeffding’s lemma,
$\mathbb{E}\!\left[\exp\!\left(\lambda\,\langle g_{t}-\mathbb{E}[g_{t}\mid\mathcal{F}_{t-1}],u\rangle\right)\Bigm|\mathcal{F}_{t-1}\right]\leq\exp\!\left(\frac{\lambda^{2}\,4G^{2}\,\|u\|^{2}}{2}\right),$
establishing conditional sub-Gaussianity.
Before quantifying the dependence of $\tau_{\varepsilon}$ on the target accuracy $\varepsilon$, it is convenient to reduce the stopping criterion $U_{t}^{\mathrm{obs}}(\alpha)\leq\varepsilon$ to a simple threshold condition on the accumulated stepsizes $S_{t}$. The following corollary makes this reduction precise.
Corollary 3.3 (General certified stopping bound).
Let Assumptions 3.2 and 1.2 hold, and let $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ be the anytime-valid upper confidence sequence from Corollary 2.4. Moreover, define
$K_{\alpha}:=\frac{1}{2}\Bigg(C\sqrt{\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log\!\big(\mathrm{e}+\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}\big)\Big)}+R_{x}^{2}+G^{2}V_{\infty}\Bigg),$ (34)
with constant $C$ defined in (10) and $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}:=\max\{4G^{2}R_{x}^{2}V_{\infty},2(\log\tfrac{2}{\alpha}+1)\}$. Then, for every target accuracy $\varepsilon>0$,
$\tau_{\varepsilon}\;\leq\;\inf\!\left\{t\geq 1:\;S_{t}\geq\frac{K_{\alpha}}{\varepsilon}\right\}\qquad\text{almost surely}.$
Corollary 3.3 shows that the stopping time $\tau_{\varepsilon}$ is controlled entirely by the growth of $S_{t}$: any deterministic lower bound on $S_{t}$ immediately yields an almost-sure upper bound on $\tau_{\varepsilon}$. The constant $K_{\alpha}$ depends only on the global problem parameters $(G,R_{x})$, the confidence level $\alpha$, and the stepsize schedule $\{\eta_{t}\}_{t\geq 1}$. In what follows, we apply Corollary 3.3 to several standard choices of stepsizes to obtain certified stopping-time complexity guarantees.
(1) Polynomially decaying stepsizes. Consider
$\eta_{t}=\eta_{0}t^{-\gamma},\qquad\eta_{0}>0,\ \ \gamma\in\Big(\frac{1}{2},1\Big).$ (35)
Since $\gamma>1/2$, this family satisfies $\sum_{t=1}^{\infty}\eta_{t}=\infty$ and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$. Lemma A.4 implies that for all $t\geq 1$,
$S_{t}\geq\frac{\eta_{0}}{1-\gamma}\Big(t^{1-\gamma}-1\Big).$
Using Corollary 3.3, it suffices that
$\frac{\eta_{0}}{1-\gamma}\Big(t^{1-\gamma}-1\Big)\ \geq\ \frac{K_{\alpha}}{\varepsilon}.$
Solving for $t$ yields:
$\tau_{\varepsilon}\ \leq\ \Biggl(1+\frac{(1-\gamma)K_{\alpha}}{\eta_{0}\,\varepsilon}\Biggr)^{\!\!\frac{1}{1-\gamma}}\qquad\text{almost surely}.$ (36)
We summarize this in the following proposition.
Proposition 3.4 (Certified stopping complexity for $\gamma\in(1/2,1)$).
Under the conditions of Corollary 3.3 and for the stepsize schedule (35), the stopping time $\tau_{\varepsilon}$ defined in (32) satisfies
$\tau_{\varepsilon}=O\!\Big(\varepsilon^{-\frac{1}{1-\gamma}}\Big)\qquad\text{almost surely}.$
In particular, $\tau_{\varepsilon}$ scales polynomially in $1/\varepsilon$ with an exponent determined by the decay rate $\gamma$.
(2) Harmonic stepsizes. Now consider the boundary case
$\eta_{t}=\frac{\eta_{0}}{t},\qquad\eta_{0}>0.$ (37)
Then $\sum_{t=1}^{\infty}\eta_{t}=\infty$ and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$, so Assumption 1.2 holds. Lemma A.5 gives $S_{t}\geq\eta_{0}\log t$. By Corollary 3.3, it suffices that
$\eta_{0}\log t\ \geq\ \frac{K_{\alpha}}{\varepsilon},$
which implies
@@ 3.1 Stopping-time complexity
$\tau_{\varepsilon}\ \leq\ \exp\!\Big(\frac{K_{\alpha}}{\eta_{0}\,\varepsilon}\Big)\qquad\text{almost surely}.$ (38)
We summarize this in the following proposition.
Proposition 3.5 (Certified stopping complexity for $\gamma=1$).
Under the conditions of Corollary 3.3 and for the stepsize schedule (37), the stopping time $\tau_{\varepsilon}$ defined in (32) satisfies
$\tau_{\varepsilon}=\exp\!\Big(O(\varepsilon^{-1})\Big)\qquad\text{almost surely}.$
Thus harmonic stepsizes yield certified stopping but with very slow complexity in $1/\varepsilon$.
Remark 3.6 (Practical implications).
The polynomial regime $\gamma\in(1/2,1)$ produces certified complexity $\tau_{\varepsilon}=O\Big(\varepsilon^{-1/(1-\gamma)}\Big)$, which matches the classical stochastic approximation tradeoff between stepsize decay and variance control. The harmonic schedule $\eta_{t}=\eta_{0}/t$, although valid for confidence-sequence stopping, leads to exponential complexity in $1/\varepsilon$ and can therefore be unattractive in practice. Combined with Theorem 3.1, these results show that the stepsize schedule entirely governs the certified stopping-time rate (up to problem-dependent constants), and that the confidence-sequence approach preserves standard stochastic approximation complexity guarantees while remaining valid under arbitrary, data-dependent optional stopping. $\clubsuit$
Acknowledgements
Appendix A Technical lemmas
Lemma A.1.
Let Assumptions 1.2 and  1.1(ii) hold. Then,
@@ Appendix A Technical lemmas
Define the nonnegative partial sums
$Y_{n}:=\sum_{t=1}^{n}\eta_{t}^{2}\|g_{t}\|^{2}\qquad\forall n\geq 1,$
and let $Y:=\sum_{t=1}^{\infty}\eta_{t}^{2}\|g_{t}\|^{2}:=\lim_{n\to\infty}Y_{n}$ (possibly taking values in $0,\infty]$). Since the sequence $\{Y_{n}\}_{n\geq 1}$ is nondecreasing and nonnegative, we may apply the monotone convergence theorem:
$\mathbb{E}[Y]=\mathbb{E}\Big[\lim_{n\to\infty}Y_{n}\Big]=\lim_{n\to\infty}\mathbb{E}[Y_{n}].$
For each $n$, we have
$\mathbb{E}[Y_{n}]=\sum_{t=1}^{n}\eta_{t}^{2}\,\mathbb{E}[\|g_{t}\|^{2}].$
Using the tower property and Assumption 1.1(ii),
$\mathbb{E}[\|g_{t}\|^{2}]=\mathbb{E}\big[\mathbb{E}[\|g_{t}\|^{2}\mid\mathcal{F}_{t-1}]\big]\leq\mathbb{E}[G^{2}]=G^{2},$
so that
$\mathbb{E}[Y_{n}]\leq\sum_{t=1}^{n}\eta_{t}^{2}G^{2}=G^{2}\sum_{t=1}^{n}\eta_{t}^{2}.$
Letting $n\to\infty$ and using Assumption 1.2, we obtain
$\mathbb{E}[Y]=\lim_{n\to\infty}\mathbb{E}[Y_{n}]\leq G^{2}\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty.$
Thus $Y$ is a nonnegative random variable with finite expectation. If $\mathbb{P}(Y=\infty)>0$, we would have $\mathbb{E}[Y]=\infty$, a contradiction. Hence
$\mathbb{P}\!\left(\sum_{t=1}^{\infty}\eta_{t}^{2}\|g_{t}\|^{2}<\infty\right)=1,$
which completes the proof. ∎
Lemma A.2.
Let $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}>1$ and define
$c_{1}\;:=\;\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1\Big)}{\log\Big(\log(\mathrm{e}+1)\Big)}\right\}.$
Fix any $L_{\alpha}\geq 1$. Then, for all $v$ with $L_{\alpha}<v\leq\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$,
$\log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;c_{1}\,\log\Big(\log(v+\mathrm{e})\Big).$
@@
First note that for $v>L_{\alpha}/\mathrm{e}$ we have $v/L_{\alpha}>1/\mathrm{e}$, hence $\log(v/L_{\alpha})>-1$, so $\log(v/L_{\alpha})+1>0$ and $\log(\log(v/L_{\alpha})+1)$ is well-defined. Moreover, $v+\mathrm{e}>\mathrm{e}$ implies $\log(v+\mathrm{e})>1$ and thus $\log(\log(v+\mathrm{e}))>0$. We now prove the inequality. Since $L_{\alpha}\geq 1$ and the logarithm is an increasing function,
$L_{\alpha}\leq v\leq\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}\ \implies\ \log(v/L_{\alpha})+1\leq\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1\ \implies\ \log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;\log(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1).$
Similarly, since $v\geq L_{\alpha}\geq 1$, we have
$v\geq 1\ \implies\ \log(v+\mathrm{e})\geq\log(\mathrm{e}+1)>1\ \implies\ \log(\log(v+\mathrm{e}))\geq\log(\log(\mathrm{e}+1)).$
Now, by the definition of $c_{1}$,
$c_{1}\;\geq\;\frac{\log(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1)}{\log(\log(\mathrm{e}+1))}\ \implies\ \log(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1)\;\leq\;c_{1}\,\log(\log(\mathrm{e}+1)).$
Combining the last three displays, we obtain for all $L_{\alpha}\leq v\leq\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$,
$\log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;\log(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1)\;\leq\;c_{1}\,\log(\log(\mathrm{e}+1))\;\leq\;c_{1}\,\log(\log(v+\mathrm{e})),$
which is the desired inequality. ∎
Lemma A.3.
Fix any $L_{\alpha}\geq 1$ and define $\nu_{0}:=2(L_{\alpha}+1)$. Then, for all $v\geq\nu_{0}$,

### v4

@@ Appendix A Additional proofs
Thus $\mathcal{E}_{\alpha}\subseteq\{f(\bar{x}_{\tau_{\varepsilon}})-f(x^{\star})\leq\varepsilon\}$ and taking probabilities yields
$\mathbb{P}\Big(f(\bar{x}_{\tau_{\varepsilon}})-f(x^{\star})\leq\varepsilon\Big)\;\geq\;\mathbb{P}(\mathcal{E}_{\alpha})\;\geq\;1-\alpha.$
This establishes (13) and completes the proof.
A.5 Proof of Lemma 3.4
Fix an outcome $\omega$ in the almost sure event on which (ii) holds for all $t\geq 1$. If $t\geq T(\varepsilon/K(\omega))$, then by definition of $T$ and the monotonicity of $b$ we have $b(t)\leq\varepsilon/K(\omega)$, and thus $U_{t}^{\mathrm{obs}}(\alpha)(\omega)\leq K(\omega)b(t)\leq\varepsilon$. By definition of $\tau_{\varepsilon}$ as the first time the certificate falls below $\varepsilon$, this implies $\tau_{\varepsilon}(\omega)\leq T(\varepsilon/K(\omega))$. Taking expectations yields the claim.
A.6 Proof of Theorem 3.6
We bound the stopping time by constructing an almost sure envelope of the form $U_{t}^{\mathrm{obs}}(\alpha)\leq K_{1}\,b(t)$ with $b(t)\asymp(1+\log t)/\sqrt{t}$ and a random prefactor $K_{1}$ having sufficiently high moments, then invoke Lemma 3.4. Recall the decomposition (as in the proof of Proposition 2.10)
$U_{t}^{\mathrm{obs}}(\alpha)=\frac{1}{2S_{t}}\Big(A_{t}(\alpha)+R_{x}^{2}+B_{t}\Big),\qquad B_{t}=\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2},$
and the definitions of $A_{t}(\alpha)$ and $\widetilde{\Sigma}_{t,\mathrm{eff}}^{2}$ used there. For the square-root schedule $\eta_{s}=\eta_{0}\,s^{-1/2}$, the deterministic estimates (39) and (40) from Proposition 2.10 yield
$S_{t}\geq\eta_{0}\sqrt{t},\qquad V_{t}\leq\eta_{0}^{2}(1+\log t).$ (42)
Moreover, the bound on the time-uniform term $A_{t}(\alpha)$ established in the same proof implies that there exists $C_{A}<\infty$ such that
$A_{t}(\alpha)\leq C_{A}(1+\log t)\qquad\text{for all }t\geq 1.$ (43)
We next control $B_{t}$ pathwise. Writing $g_{s}=\nabla f(x_{s})+\xi_{s}$ and using $\|a+b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$ together with Assumption 3.3, we obtain
$B_{t}=\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\leq 2\sum_{s=1}^{t}\eta_{s}^{2}\|\nabla f(x_{s})\|^{2}+2\sum_{s=1}^{t}\eta_{s}^{2}\|\xi_{s}\|^{2}\leq 2G^{2}V_{t}+2\sum_{s=1}^{t}\eta_{s}^{2}\|\xi_{s}\|^{2}.$
Decompose the weighted noise-square sum into its predictable compensator plus a martingale:
$\sum_{s=1}^{t}\eta_{s}^{2}\|\xi_{s}\|^{2}=\sum_{s=1}^{t}\eta_{s}^{2}\,\mathbb{E}[\|\xi_{s}\|^{2}\mid\mathcal{F}_{s-1}]+M_{t},\qquad M_{t}:=\sum_{s=1}^{t}\Delta_{s},$
where $\Delta_{s}:=\eta_{s}^{2}(\|\xi_{s}\|^{2}-\mathbb{E}[\|\xi_{s}\|^{2}\mid\mathcal{F}_{s-1}])$. By Lemma B.2, $\mathbb{E}[\|\xi_{s}\|^{2}\mid\mathcal{F}_{s-1}]\leq 4d\sigma^{2}$ almost surely, hence
$\sum_{s=1}^{t}\eta_{s}^{2}\,\mathbb{E}[\|\xi_{s}\|^{2}\mid\mathcal{F}_{s-1}]\leq 4d\sigma^{2}V_{t}\leq 4d\sigma^{2}\eta_{0}^{2}(1+\log t).$
Combining the last two displays and using $V_{t}\leq\eta_{0}^{2}(1+\log t)$ again gives that there exists $C_{B}<\infty$ (depending only on $(G,\sigma,d,\eta_{0})$) such that
$B_{t}\leq C_{B}(1+\log t)+2\sup_{s\geq 1}|M_{s}|.$ (44)
We now introduce the random variable $K:=1+\sup_{s\geq 1}|M_{s}|$. By Lemma B.3, $\mathbb{E}[K^{4}]<\infty$, and in particular $K<\infty$ almost surely. Using (43), (44), and $S_{t}\geq\eta_{0}\sqrt{t}$, we obtain that there exists $C_{0}<\infty$ depending only on $(\alpha,\eta_{0},R_{x},G,\sigma,d)$ such that for all $t\geq 1$,
@@
$U_{t}^{\mathrm{obs}}(\alpha)\leq C_{0}\,K\,\frac{1+\log t}{\sqrt{t}}\qquad\text{almost surely.}$ (45)
Define $b(t):=(1+\log t)/\sqrt{t}$ for $t\geq 1$. Then $b(t)\to 0$ as $t\to\infty$, and $b$ is nonincreasing for all sufficiently large $t$. Recall the definition $T(u):=\inf\{t\geq 1:\ b(t)\leq u\}$. As shown in Lemma B.7, there exists a universal constant $C_{1}<\infty$ such that for all $u\in(0,1/2)$,
$T(u)\leq C_{1}\,u^{-2}\,\log^{2}\!\frac{1}{u}.$ (46)
Set $K_{1}:=C_{0}K\geq 1$. From (45) we have $U_{t}^{\mathrm{obs}}(\alpha)\leq K_{1}\,b(t)$ almost surely, and therefore Lemma 3.4 yields
$\mathbb{E}[\tau_{\varepsilon}]\leq\mathbb{E}\!\left[T\!\left(\frac{\varepsilon}{K_{1}}\right)\right].$
Using (46) and $\varepsilon\in(0,1/2)$ gives
$\mathbb{E}[\tau_{\varepsilon}]\leq C_{1}\,\varepsilon^{-2}\,\mathbb{E}\!\left[K_{1}^{2}\,\log^{2}\!\left(\frac{K_{1}}{\varepsilon}\right)\right].$
To bound the expectation uniformly in $\varepsilon$, note that $\log(K_{1}/\varepsilon)\leq\log(1/\varepsilon)+\log K_{1}$, so
$K_{1}^{2}\log^{2}\!\left(\frac{K_{1}}{\varepsilon}\right)\leq 2K_{1}^{2}\log^{2}\!\frac{1}{\varepsilon}+2K_{1}^{2}\log^{2}(K_{1}).$
The first term has expectation $2\log^{2}(1/\varepsilon)\,\mathbb{E}[K_{1}^{2}]<\infty$.
For the second term, fix any $\delta\in(0,1)$. Since $x^{\delta}$ dominates $\log^{2}x$ as $x\to\infty$ and $x\mapsto\log^{2}x/x^{\delta}$ is continuous on $1,\infty)$, there exists $C_{\delta}<\infty$ such that $\log^{2}x\leq C_{\delta}x^{\delta}$ for all $x\geq 1$. Because $K_{1}\geq 1$ almost surely, this yields $K_{1}^{2}\log^{2}(K_{1})\leq C_{\delta}K_{1}^{2+\delta}$ almost surely. Because $\mathbb{E}[K_{1}^{2+\delta}]<\infty$ (which holds since $\mathbb{E}[K^{4}]<\infty$ and $\delta\leq 2$), it follows that $\mathbb{E}[K_{1}^{2}\log^{2}(K_{1})]<\infty$. Therefore there exists a constant $C_{2}<\infty$, independent of $\varepsilon$, such that
$\mathbb{E}\!\left[K_{1}^{2}\,\log^{2}\!\left(\frac{K_{1}}{\varepsilon}\right)\right]\leq C_{2}\,\log^{2}\!\frac{1}{\varepsilon}.$
Combining the last two displays yields
$\mathbb{E}[\tau_{\varepsilon}]\leq C\,\varepsilon^{-2}\,\log^{2}\!\Big(\frac{1}{\varepsilon}\Big),$
after absorbing constants into $C$, which concludes the proof.
Appendix B Supporting lemmas
Lemma B.1.
Let Assumption 2.7 hold and assume that $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$. Then,
@@ Appendix B Supporting lemmas
Let $Y_{n}:=\sum_{t=1}^{n}\eta_{t}^{2}\|g_{t}\|^{2}$ and $Y:=\lim_{n\to\infty}Y_{n}\in[0,\infty]$. By monotone convergence,
$\mathbb{E}[Y]=\lim_{n\to\infty}\mathbb{E}[Y_{n}]=\lim_{n\to\infty}\sum_{t=1}^{n}\eta_{t}^{2}\,\mathbb{E}[\|g_{t}\|^{2}].$
By the tower property and Assumption 2.7, $\mathbb{E}[\|g_{t}\|^{2}]=\mathbb{E}[\mathbb{E}[\|g_{t}\|^{2}\mid\mathcal{F}_{t-1}]]\leq M^{2}$. Hence
$\mathbb{E}[Y]\leq M^{2}\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty.$
Since $Y\geq 0$, $\mathbb{E}[Y]<\infty$ implies $\mathbb{P}(Y<\infty)=1$. ∎
Lemma B.2.
Suppose Assumption 1.2 holds. Then, for every $t\geq 1$, [^1] The constants in this lemma are not optimized. Under (2), a sharper moment computation yields $\mathbb{E}[\|\xi_{t}\|^{2}\mid\mathcal{F}_{t-1}]\leq d\sigma^{2}$, $\mathbb{E}[\|\xi_{t}\|^{4}\mid\mathcal{F}_{t-1}]\leq 3d^{2}\sigma^{4}$, and $\mathbb{E}[\|\xi_{t}\|^{8}\mid\mathcal{F}_{t-1}]\leq 105\,d^{4}\sigma^{8}$. We keep the present bounds for simplicity.
$\mathbb{E}[\|\xi_{t}\|^{2}\mid\mathcal{F}_{t-1}]\leq 4d\sigma^{2},\quad\mathbb{E}[\|\xi_{t}\|^{4}\mid\mathcal{F}_{t-1}]\leq 16d^{2}\sigma^{4},\quad\mathbb{E}[\|\xi_{t}\|^{8}\mid\mathcal{F}_{t-1}]\leq 768\,d^{4}\sigma^{8}\qquad\text{almost surely.}$
@@
Fix $t\geq 1$ and $i\in\{1,\dots,d\}$. Taking $u=e_{i}$ in (2) yields
$\mathbb{E}\!\left[e^{\lambda\xi_{t,i}}\mid\mathcal{F}_{t-1}\right]\leq\exp\!\left(\frac{\lambda^{2}\sigma^{2}}{2}\right)\qquad\forall\lambda\in\mathbb{R}.$
By a standard Chernoff bound (see [Boucheron et al., 2013, Section 2.3]), this implies
$\mathbb{P}(\xi_{t,i}>x\mid\mathcal{F}_{t-1})\ \vee\ \mathbb{P}(-\xi_{t,i}>x\mid\mathcal{F}_{t-1})\leq\exp\!\left(-\frac{x^{2}}{2\sigma^{2}}\right),\qquad\forall x>0.$
Thus, conditional on $\mathcal{F}_{t-1}$, $\xi_{t,i}$ satisfies the assumption in [Boucheron et al., 2013, Theorem 2.1] with $v=\sigma^{2}$. Applying that theorem with $q=1$, $q=2$, and $q=4$ yields
$\mathbb{E}[\xi_{t,i}^{2}\mid\mathcal{F}_{t-1}]\leq 4\sigma^{2},\qquad\mathbb{E}[\xi_{t,i}^{4}\mid\mathcal{F}_{t-1}]\leq 16\sigma^{4},\qquad\mathbb{E}[\xi_{t,i}^{8}\mid\mathcal{F}_{t-1}]\leq 768\,\sigma^{8},$
almost surely. Summing the $q=1$ inequality over $i$ gives
$\mathbb{E}[\|\xi_{t}\|^{2}\mid\mathcal{F}_{t-1}]=\sum_{i=1}^{d}\mathbb{E}[\xi_{t,i}^{2}\mid\mathcal{F}_{t-1}]\leq 4d\sigma^{2}.$
Next, using $(\sum_{i=1}^{d}a_{i})^{2}\leq d\sum_{i=1}^{d}a_{i}^{2}$ with $a_{i}=\xi_{t,i}^{2}\geq 0$,
$\|\xi_{t}\|^{4}=\Big(\sum_{i=1}^{d}\xi_{t,i}^{2}\Big)^{2}\leq d\sum_{i=1}^{d}\xi_{t,i}^{4},$
and hence
$\mathbb{E}[\|\xi_{t}\|^{4}\mid\mathcal{F}_{t-1}]\leq d\sum_{i=1}^{d}\mathbb{E}[\xi_{t,i}^{4}\mid\mathcal{F}_{t-1}]\leq 16d^{2}\sigma^{4}.$
Finally, using $(\sum_{i=1}^{d}a_{i})^{4}\leq d^{3}\sum_{i=1}^{d}a_{i}^{4}$ with $a_{i}=\xi_{t,i}^{2}\geq 0$,
$\|\xi_{t}\|^{8}=\Big(\sum_{i=1}^{d}\xi_{t,i}^{2}\Big)^{4}\leq d^{3}\sum_{i=1}^{d}\xi_{t,i}^{8},$
and therefore
$\mathbb{E}[\|\xi_{t}\|^{8}\mid\mathcal{F}_{t-1}]\leq d^{3}\sum_{i=1}^{d}\mathbb{E}[\xi_{t,i}^{8}\mid\mathcal{F}_{t-1}]\leq d^{3}d\,768\,\sigma^{8}=768\,d^{4}\sigma^{8}.$
This concludes the proof. ∎
Lemma B.3.
Suppose Assumption 1.2 holds. Let $\{\eta_{t}\}_{t\geq 1}$ be a deterministic stepsize sequence satisfying $\sum_{t=1}^{\infty}\eta_{t}^{4}<\infty$, and define the adapted process $\{M_{t}\}_{t\geq 1}$ by
$M_{t}:=\sum_{s=1}^{t}\Delta_{s},\qquad\Delta_{s}:=\eta_{s}^{2}\Big(\|\xi_{s}\|^{2}-\mathbb{E}[\|\xi_{s}\|^{2}\mid\mathcal{F}_{s-1}]\Big).$
Then $\{M_{t}\}$ is an $\{\mathcal{F}_{t}\}$-martingale and
$\mathbb{E}\!\left[\sup_{t\geq 1}|M_{t}|^{4}\right]<\infty.$
Proof.
Since $\Delta_{s}$ is $\mathcal{F}_{s}$-measurable, integrable, and satisfies $\mathbb{E}[\Delta_{s}\mid\mathcal{F}_{s-1}]=0$ almost surely by construction, it follows that $\{M_{t}\}_{t\geq 1}$ is an $\{\mathcal{F}_{t}\}$-martingale with differences $\{\Delta_{s}\}_{s\geq 1}$. Fix $T\geq 1$. By Doob’s $L^{4}$ maximal inequality for martingales (e.g., [Durrett, 2019, Theorem 4.4.4]),
$\mathbb{E}\!\left[\sup_{1\leq t\leq T}|M_{t}|^{4}\right]\leq\Big(\frac{4}{3}\Big)^{4}\,\mathbb{E}\!\left[|M_{T}|^{4}\right].$ (47)
Next, Burkholder’s square-function inequality (upper bound in [Hall and Heyde, 2014, Theorem 2.10]) applied to the martingale $\{M_{t}\}_{t=0}^{T}$ with differences $\Delta_{s}$ yields an (explicit) constant $C_{1}<\infty$ such that
$\mathbb{E}\!\left[|M_{T}|^{4}\right]\leq C_{1}\,\mathbb{E}\!\left[\Big(\sum_{s=1}^{T}\Delta_{s}^{2}\Big)^{2}\right].$ (48)
Combining (47) and (48) gives
$\mathbb{E}\!\left[\sup_{1\leq t\leq T}|M_{t}|^{4}\right]\leq\Big(\frac{4}{3}\Big)^{4}C_{1}\,\mathbb{E}\!\left[\Big(\sum_{s=1}^{T}\Delta_{s}^{2}\Big)^{2}\right].$ (49)
Since $\Delta_{s}^{2}\geq 0$, Minkowski’s inequality in $L^{2}$ implies
$\Big\|\sum_{s=1}^{T}\Delta_{s}^{2}\Big\|_{L^{2}}\leq\sum_{s=1}^{T}\|\Delta_{s}^{2}\|_{L^{2}}=\sum_{s=1}^{T}\|\Delta_{s}\|_{L^{4}}^{2},$
where $\|\Delta_{s}^{2}\|_{L^{2}}=(\mathbb{E}|\Delta_{s}|^{4})^{1/2}=\|\Delta_{s}\|_{L^{4}}^{2}$. Therefore,
$\mathbb{E}\!\left[\Big(\sum_{s=1}^{T}\Delta_{s}^{2}\Big)^{2}\right]=\Big\|\sum_{s=1}^{T}\Delta_{s}^{2}\Big\|_{L^{2}}^{2}\leq\Big(\sum_{s=1}^{T}(\mathbb{E}|\Delta_{s}|^{4})^{1/2}\Big)^{2}.$ (50)
To bound $\mathbb{E}|\Delta_{s}|^{4}$, set $A_{s}:=\|\xi_{s}\|^{2}\geq 0$ and $\mu_{s}:=\mathbb{E}[A_{s}\mid\mathcal{F}_{s-1}]$. The inequality $|a-b|^{4}\leq 8(a^{4}+b^{4})$ gives $|A_{s}-\mu_{s}|^{4}\leq 8(A_{s}^{4}+\mu_{s}^{4})$, and taking conditional expectations yields
$\mathbb{E}\!\left[|A_{s}-\mu_{s}|^{4}\,\middle|\,\mathcal{F}_{s-1}\right]\leq 8\,\mathbb{E}[A_{s}^{4}\mid\mathcal{F}_{s-1}]+8\,\mu_{s}^{4}.$
By conditional Jensen’s inequality, $\mu_{s}^{4}=(\mathbb{E}[A_{s}\mid\mathcal{F}_{s-1}])^{4}\leq\mathbb{E}[A_{s}^{4}\mid\mathcal{F}_{s-1}]$, and hence
$\mathbb{E}\!\left[|A_{s}-\mu_{s}|^{4}\,\middle|\,\mathcal{F}_{s-1}\right]\leq 16\,\mathbb{E}[A_{s}^{4}\mid\mathcal{F}_{s-1}]=16\,\mathbb{E}\!\left[\|\xi_{s}\|^{8}\mid\mathcal{F}_{s-1}\right].$
Consequently,
$\mathbb{E}\!\left[|\Delta_{s}|^{4}\mid\mathcal{F}_{s-1}\right]=\eta_{s}^{8}\,\mathbb{E}\!\left[|A_{s}-\mu_{s}|^{4}\mid\mathcal{F}_{s-1}\right]\leq 16\,\eta_{s}^{8}\,\mathbb{E}\!\left[\|\xi_{s}\|^{8}\mid\mathcal{F}_{s-1}\right].$
Under Assumption 1.2 (conditional sub-Gaussian noise), the conditional $8$th moment is uniformly bounded; in particular, there exists $C_{2}<\infty$ such that $\mathbb{E}[\|\xi_{s}\|^{8}\mid\mathcal{F}_{s-1}]\leq C_{2}$ almost surely for all $s$ (e.g., Lemma B.2). Taking expectations gives
$\mathbb{E}|\Delta_{s}|^{4}\leq 16C_{2}\,\eta_{s}^{8},\qquad\text{so}\qquad(\mathbb{E}|\Delta_{s}|^{4})^{1/2}\leq(16C_{2})^{1/2}\eta_{s}^{4}.$
Plugging this bound into (50) yields
$\mathbb{E}\!\left[\Big(\sum_{s=1}^{T}\Delta_{s}^{2}\Big)^{2}\right]\leq(16C_{2})\Big(\sum_{s=1}^{T}\eta_{s}^{4}\Big)^{2}\leq(16C_{2})\Big(\sum_{s=1}^{\infty}\eta_{s}^{4}\Big)^{2}<\infty.$
Combining with (49) shows that there exists a deterministic constant $C_{3}<\infty$ such that
$\sup_{T\geq 1}\mathbb{E}[\sup_{1\leq t\leq T}|M_{t}|^{4}]\leq C_{3}.$
Since $\sup_{1\leq t\leq T}|M_{t}|^{4}$ increases to $\sup_{t\geq 1}|M_{t}|^{4}$ as $T\to\infty$, monotone convergence yields
$\mathbb{E}\!\left[\sup_{t\geq 1}|M_{t}|^{4}\right]\leq C_{3}<\infty,$
which concludes the proof. ∎
Lemma B.4.
Fix any $L_{\alpha}\geq 1$ and define $v_{0}:=2(L_{\alpha}+1)$. Then, for all $v\geq v_{0}$,